OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

jimthompson5910 (jim_thompson5910):

use the point-slope formula y - y1 = m(x-x1) where m = slope x1 = x coordinate of given point y1 = y coordinate of given point

OpenStudy (anonymous):

ohhh okkk so is it y + 2 = 6(x + 3) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

very close, but the slope is negative, not positive

jimthompson5910 (jim_thompson5910):

oh i don't know why i didn't see that the terms are flipped

jimthompson5910 (jim_thompson5910):

my bad lol

jimthompson5910 (jim_thompson5910):

m = -6 x1 = -2 y1 = -3 y - y1 = m(x-x1) y - (-3) = -6(x-(-2)) y+3 = -6(x+2)

OpenStudy (anonymous):

@jim_thompson5910 thank you so much!! :) i get it now :))

jimthompson5910 (jim_thompson5910):

yeah there's not much to it yw